How many gallons of a 20% ammonia solution must be mixed with a 5% ammonia solution to create 20 gallons of 11% ammonia solution?
Answer by stanbon(48558)

How many gallons of a 20% ammonia solution must be mixed with a 5% ammonia solution to create 20 gallons of 11% ammonia solution?
Let "x" be amount of 20% solution; amount of ammonia is 0.2x gallons
The "20-x" is amount of 5% solution: amount of ammonia is
0.05(20-x)=0.1-0.02x gallons
Amount of mixture is 20 gallons: amount of ammonia in it is 0.11(20)=2.2 gallons
-----------
EQUATION:
ammonia + ammonia = ammonia
0.2x + 0.1-0.02x=2.2
Multiply thru by 100 to get:
20x+10-2x=220
18x=210
x= 11 2/3 gallons (amount of 20% solution)
20-x= 8 1/3 gallons (amount of 5% solution)
Cheers,
Stan H.